### v3

# Directional Sign Loss: A Topology-Preserving Loss Function that Approximates the Sign of Finite Differences

Abstract
Preserving topological features in learned latent spaces is a fundamental challenge in representation learning, particularly for topology-sensitive data. This paper introduces directional sign loss (DSL), an efficient, differentiable loss function that approximates the number of mismatches in the signs of finite differences between corresponding elements of two arrays. By penalizing discrepancies in critical points between input and reconstructed data, DSL encourages autoencoders and other learnable compressors to retain the topological features of the original data. We present the formulation and complexity analysis of DSL, comparing it to other non-differentiable topological measures. Experiments on multidimensional array data show that combining DSL with traditional loss functions preserves topological features more effectively than traditional losses alone. DSL serves as a differentiable, efficient proxy for common topology-based metrics, enabling topological feature preservation on previously impractical problem sizes and in a wider range of gradient-based optimization frameworks.
1 Introduction
Learned representations are central to modern machine learning, serving as compressed yet useful descriptions of data. While representations are often optimized for specific tasks, reconstructions are equally important, not only for downstream machine learning tasks, but also for human interpretability and applications. Standard loss functions consider only point-wise similarity and often fail to capture topological structures in data. This can lead to reconstructions that appear faithful but miss geometric or structural features. While topology-aware learning methods, such as persistent homology, offer approaches to capturing these features, they are typically computationally inefficient and not differentiable, preventing their incorporation into gradient-based learning frameworks with large datasets.
Topological features capture the innate structure of data. These features include connected components, loops, and critical points (maxima, minima, and saddle points) (Carlsson, 2009) and are often obtained using topological data analysis (TDA) tools such as persistence diagrams. These features can be used for analyzing and visualizing the data structure, specifically for data where topology plays a vital role. Distance measures such as the Wasserstein distance describe the dissimilarity between the topological features of two pieces of data (Edelsbrunner and Harer, 2010). These measures help researchers draw out relevant information while filtering out noise or irrelevant variations in the data and are useful in machine learning (ML) tasks such as classifying or clustering topology-rich data (Maljovec et al., 2013; Yan et al., 2021). Scientific simulations, including those in fluid dynamics, physics, and weather prediction, leverage topological features to identify phenomena such as vortices, changes in atomic formation, and pressure extrema (Agarwal et al., 2021; Ramamurthi et al., 2022). Similarly, medical imaging often relies on capturing and comparing topological structures, such as gene interactions and tumors (Singh et al., 2023). Economists also examine financial networks through persistent features to detect market anomalies and gain insights into overall market behavior (Mojdehi et al., 2025; Eberle and Holder, 2007). Preservation of topological invariants is often essential for reliable interpretation and downstream decision-making, even at the expense of numerical accuracy (Tierny, 2017).
Lossy data compressors can reduce storage and data transfer requirements for ML models trained on large datasets, aiding data preprocessing, training, and inference. Compressed data also lowers memory bandwidth usage, speeding GPU or TPU operations and accelerating experiments. However, these compressors, including learnable variants, typically minimize point-wise error rather than preserve topological properties (Soler et al., 2018a). As a result, they can distort critical points or introduce new ones (Diffenderfer et al., 2019), propagating errors through the data processing pipeline and leading to inaccurate outputs.
To address the interests and motivations of both topological data analysis and lossy data compression, we introduce a novel loss function that approximates the number of mismatches in the signs of finite differences between the corresponding elements of input and reconstructed data.
The key contributions in the paper are as follows:
  • Directional sign loss (DSL), a differentiable measure of the dissimilarity between the critical points of two arrays.
  • An analysis of its time and memory complexity, as well as a demonstration of its behavior on a simple example.
  • Evaluations of learnable compressors that show that this loss function can be used as a proxy for other measures of topological dissimilarity, consisting of two simple autoencoders and one variational autoencoder, using this loss function, trained on 1-dimensional arrays of stock prices, 2-dimensional images, and 3-dimensional physics simulation states.
In the rest of this paper, we introduce the background and related work (§ 2.2), describe the loss function and demonstrate its behavior on a simple example (§ 3), evaluate it by training several autoencoders and discussing the results (§ 4), discuss limitations (§ 5), and offer our conclusion (§ 6).
2 Background and Related Work
We describe the background on autoencoders and topological data analysis, and discuss related work.
2.1 Background
2.1.1 Autoencoders
Autoencoders (AEs) (Hinton and Salakhutdinov, 2006) are neural networks that learn efficient representations of data by training on a reconstruction task. They consist of an encoder which transforms inputs to a lower-dimensional compressed representation, and a decoder which reconstructs the original input from the compressed representation. By optimizing them using a loss function that penalizes deviation from the original, AEs learn to capture essential features while discarding irrelevant information with respect to the loss function. Variants of AEs such as variational autoencoders (VAEs) (Kingma and Welling, 2022), adversarial autoencoders (Makhzani et al., 2016), denoising autoencoders (Vincent et al., 2008), and sparse autoencoders (Schölkopf et al., 2007) introduce additional features such as generative modeling, robustness, and interpretability.
2.1.2 Topological data analysis
Critical points: A critical point of a continuous function $f$ is a point in the domain where $\nabla f=0$ (Edelsbrunner and Harer, 2010). However, in discrete settings such as sampled fields on a grid, the definition of critical points from continuous functions does not directly apply. For vector fields, identifying critical points can require detecting sign changes in the determinants of vectors within simplices formed from the vector field. This process involves evaluating whether replacing a vertex with zero leads to a determinant sign change, indicating a potential critical point (Xia et al., 2024). For scalar fields (“tensors”), it is sufficient to approximate critical points by identifying locations where finite differences in some direction cross zero. If a critical point exists in an underlying continuous function $f$, then a zero crossing in the finite difference $\hat{f}(x+h)-\hat{f}(x)$ of the discrete approximation $\hat{f}$, using step size $h$, serves as an indicator of an approximate critical point with $O(h)$ error (Powell, 1981).
Persistence diagrams: A persistence diagram is a topological summary that captures the birth and death of homological features in a filtered space. Given a filtration, a nested sequence of spaces defined by a function, topological features such as connected components, loops, and voids appear and disappear at specific function values. Each feature is represented as a tuple in the persistence diagram, where the first coordinate corresponds to the function value at which the feature first appears (birth) and the second coordinate corresponds to where it disappears (death). When two features merge, the older one persists while the newer one disappears. Since changes in topology occur at critical points of the filtration function, persistence diagrams implicitly encode information about critical points. Persistence diagrams provide a compact representation of topological structure. For $d$-dimensional arrays, persistence diagrams can be constructed for homology orders up to ${d-1}$. However, the majority of persistent features in practice, particularly connected components (order 0), dominate the topological summary, with higher-order features becoming progressively less common in most real-world data. For a worked-through example, see Appendix 6.1.1.
Distance measures: Several metrics quantify the dissimilarity between persistence diagrams, which in turn reflect differences in topological structure. Our work uses these metrics to evaluate the preservation of topological features between original and reconstructed data. The Wasserstein distance (Malinovskii, 1975) between persistence diagrams is the most common measure of topological dissimilarity. The Wasserstein distance describes the cost of transforming one probability distribution into another, typically using norm order $p=1$ or $p=2$. The bottleneck distance is a special case of Wasserstein distance where $p=\infty$, capturing the largest discrepancy between the most dissimilar birth-death pairs. For persistence diagrams, these metrics compare differences between corresponding birth-death pairs in the two diagrams, implicitly reflecting variations in the underlying critical points of the filtration function. For a detailed description, see Appendix 6.1.3. The functional distortion distance between merge trees measures both structural and functional differences by optimally matching nodes in the merge trees based on their scalar function values, and penalizing unmatched or structurally divergent nodes. The distance is computed by matching nodes from one tree to the closest nodes in the other (in terms of function value), summing the minimal differences, and adding a penalty for differences in the number of critical points. This metric directly compares the topological structure of scalar fields, rather than their persistence diagrams. The pairwise linear correlation distance quantifies the dissimilarity of the spatial arrangements of feature vectors in the original and reconstructed data. All pairwise Euclidean distances between feature vectors at each spatial location are computed. Then the correlation distance is one minus the Pearson correlation coefficient between the resulting distance matrices for the two datasets. This metric captures the preservation of the overall geometric arrangement of features.
2.2 Related work
Lossy data compressors such as Di and Cappello (2016); Lindstrom (2014); Lakshminarasimhan et al. (2013); Lindstrom and Isenburg (2006) are usually developed to achieve a high compression ratio with an admissible fidelity loss. Some lossy compressors prioritize other features, such as differentiability or compressed-space operations (Agarwal et al., 2023, 2024a, 2024b). Others aim to guarantee the preservation of a controllable set of topological features during compression (Soler et al., 2018b; Yan et al., 2024; Xia et al., 2024). These compressors mainly focus on maintaining point-wise error along with high compression speed and ratio. While some explicitly target topological feature preservation, all of these compressors are “frozen” and are not designed to adapt to specific data distributions as learnable compressors are.
Autoencoders (AEs) have been widely used as learnable compression models. Traditional AEs and their common variants focus primarily on reconstructing the input while learning a latent space using a standard reconstruction loss, without explicitly considering the structural properties of the data.
To address this limitation, several studies have explored adding topological constraints into neural network training. Topological autoencoders (Moor et al., 2021) aim to preserve topological features of the input space in the latent representation. Such autoencoders use measures such as representation topology divergence (RTD) (Trofimov et al., 2023) to describe topological dissimilarity between data manifolds and latent representations. While these methods are similar to our work, they are meant to capture the topological features of datasets rather than individual examples, relying on small samples (up to size $\approx 2^{8}$) of small examples ($\approx 2^{16}$ elements) to retain tractability due to their reliance on persistence diagrams or distance matrices.
Persistence-weighted kernels (Kusano et al., 2016) and persistence landscapes (Bubenik, 2015) have been used to regularize learned representations and improve their structural retention. Topological layers have been used to integrate persistent homology into learning objectives (Brüel-Gabrielsson et al., 2020) using level set and edge-based filters. Topology-aware representational similarity analysis (tRSA) has appeared to characterize the topology of learned representations (Lin and Kriegeskorte, 2024), showing robustness to noise and individual variability. Unlike these approaches, our method requires no explicit topological descriptors or architectural modifications, such as specialized layers.
Our approach differs from these methods in that it does not impose constraints on the latent space or require modifications to the architecture of the AE. While methods such as Moor et al. (2021); Trofimov et al. (2023) enforce topological alignment between data and latent spaces through divergence minimization, our loss function directly encourages the preservation of topological features across compression and decompression of entire inputs (and not samples thereof) without accessing the latent space. This flexibility makes our approach particularly applicable in scenarios where computational efficiency and topological integrity must be balanced.
3 Directional Sign Loss
In many ML applications, one may prefer to capture directional trends and critical points. Standard loss functions such as the mean squared error (MSE) and mean absolute error (MAE) capture magnitude-based discrepancies but do not account for mismatches in directional changes across data points, e.g. finite differences. This limits their suitability in tasks where directional trends are emphasized, particularly where trends matter more than raw values.
To address this limitation, we introduce directional sign loss (DSL), a loss function that penalizes mismatches in the signs of finite differences between adjacent elements of predictions and targets. This encourages a model to align its predictions with the directional trends of the data.
3.1 Algorithm
In our notation, operations are performed on whole arrays, broadcasting where necessary, unless indexed into particular slices or elements, as is conventional in common deep learning frameworks.
Figure 1: $\tanh(sx)$ using different values of $s$ and other sign-like functions.
One way to capture the differences in critical points between two arrays of numbers (“tensors”) is to take their finite differences in all directions. If tensor $X\in\mathbb{R}^{n}$ were one-dimensional, this would amount to computing $X_{2:n}-X_{1:n-1}$, and doing the same to the other tensor, say $Y$. The signs of these finite differences can be expressed as integers using $\mathrm{sign}(x)=\{1\;\mathrm{if}\;x>0,0\;\mathrm{if}\;x=0,-1\;\mathrm{if}\;x<0$. Then, the number of mismatches between the signs of their finite differences would be aggregated, serving as a measure of the dissimilarity of their critical points. In contrast to other metrics, this focuses only on the presence and location of the critical points and not on the values that construct them. However, counting mismatches directly would not be differentiable. Thus, we use a smooth function $\tanh$, with a sharpness parameter $s$, to approximate the $\mathrm{sign}$ function.
DSL computes the total mismatch in directional trends between tensors $X$ and $Y\in\mathbb{R}^{n_{1}\times n_{2}\times\dots\times n_{d}}$, which must have the same shape. One could be a reconstruction or approximation of the other. For each dimension $i$ of the tensors from 111 to $d$, the following steps are performed.
  1. 1.
    Compute finite differences: Adjacent elements in each dimension are subtracted. This is done for the $i$th dimension for both $X$ and $Y$.
  2. 2.
    Approximate directional signs: The finite differences are scaled by a sharpness factor $s$, and the $\tanh$ function is applied, with larger values of $s$ producing a closer approximation to the $\mathrm{sign}$ function. Larger $s$ helps distinguish DSL from other loss functions such as MAE loss, but risks vanishing gradients, making the choice of $s$ a hyperparameter. Figure 1 shows $\tanh(sx)$ with different values of $s$. The user may use other $\mathrm{sign}$-like functions in DSL, such as $\mathrm{softsign}$ Glorot and Bengio (2010).
  3. 3.
    Compute mismatches: The absolute differences between the approximated signs of the finite differences are calculated and optionally divided by two. This division matches the loss function’s units to those of a non-differentiable exact version that uses the $\mathrm{sign}$ function and counts mismatches directly.
[…]
Input: Tensors $X$ and $Y$ of shape $(n_{1},n_{2},\dots,n_{d})$, sharpness $s$, optional weights $W$
Output: Loss value $\ell$
$\ell\leftarrow 0$;
for $i\leftarrow 1$ to $d$ do
   Compute finite differences:
$D_{X}\leftarrow X_{(:)*(i{-}1),2{:}n_{i},(:)*(d{-}i)}-X_{(:)*(i{-}1),1{:}(n_{i}{-}1),(:)*(d{-}i)}$
$D_{Y}\leftarrow Y_{(:)*(i{-}1),2{:}n_{i},(:)*(d{-}i)}-Y_{(:)*(i{-}1),1{:}(n_{i}{-}1),(:)*(d{-}i)}$
Find approximate sign mismatches:
$\Delta\leftarrow\frac{|\tanh(sD_{X})-\tanh(sD_{Y})|}{2}$
With optional weights $W$, accumulate mismatches:
$\ell\leftarrow\ell+(W_{i}\text{ or }1)\sum\Delta$
return $\ell$
Algorithm 1 Directional Sign Loss (DSL)
3.2 Time and memory complexity
We consider the time complexity using an idealized parallel computing model, with the availability of infinite parallel processors to differentiate the parts that are inherently parallelizable and those that are sequential bottlenecks. Such assumptions are commonly employed in GPU and high-performance computing literature because they illustrate the theoretical scalability and potential limitations of parallel implementations. If $X$ and $Y$ are $d$-dimensional arrays shaped $(n_{1},n_{2},\dots,n_{d})$, then we can achieve a time complexity of $O(\log_{2}d+\sum_{i=1}^{d}\log_{2}n_{i})$, but this would take $O(d\prod_{i=1}^{d}n_{i})$ memory (See Appendix 6.2). Our implementation instead resembles the pseudocode in Algorithm 1, where the mismatches are summed separately per dimension. This has a complexity of
$O\left(d\sum_{i=1}^{d}\log_{2}n_{i}\right)\text{time},\quad O\left(\prod_{i=1}^{d}n_{i}\right)\text{memory}.$
For comparison, MSE has a time complexity of $O(\sum_{i=1}^{d}\log_{2}n_{i})$ and the same memory complexity. Making persistence diagrams and taking the Wasserstein distance between them takes $O(\prod_{i=1}^{d}n_{i}^{3})$ time and $O(\prod_{i=1}^{d}n_{i})$ memory. Merge tree distance computation between trees with $n$ and $m$ extrema has an optimized time complexity of $O(n^{2}m^{2}(n+m))$ using memoization, with $O(nm)$ memory required for storing intermediate matching costs (Morozov et al., 2013; Yan et al., 2022). Without memoization, this takes $O(2^{n+m})$ time. Pairwise linear correlation analysis between $n$ variables with $m$ samples each requires $O(n^{2}m)$ time for full matrix computation, reducible to $O(n^{2}\log m)$ using fast Fourier transform windowing techniques (Pearson, 1895). Memory complexity ranges from $O(n^{2})$ for storing all correlations to $O(1)$ when computing on-the-fly. Figure 2 shows experimental time and memory usage of these loss functions.
(a) time 1D
(b) memory 1D
(c) time 2D
(d) memory 2D
(e) legend
Figure 2: Time and peak memory usage from 1-dimensional and 2-dimensional array inputs using one AMD Ryzen 7 7800X3D CPU and one NVIDIA RTX 3090 GPU (24GB). Points are not shown if the time per example exceeds 1 second or GPU memory is exceeded. 2D inputs sizes show the number of elements per side of a square. Persistence diagram distance was the 2-order Wasserstein distance between order-0 (the cheapest) persistence diagrams. 3D inputs are shown in Appendix Figure 8.
Implementation notes: In our implementation, we skip the main calculation for the first (batch) dimension and provide options for reduction (mean, sum, or none), resembling typical PyTorch loss functions. If the option to match the units of the exact version of DSL (using $\mathrm{sign}$ instead of $\tanh$) is turned off, we scale the return value by the number of comparisons, making it easier to use in combination with conventional loss functions such as MSE loss. As the choice of $s$ greatly affects the behavior of the loss function, we provide an approach to learn $s$ in Appendix 6.5.
3.3 Demonstration
To illustrate DSL’s behavior, we use a simple example. We trained 2-layer perceptrons with $\mathrm{ReLU}$ between the layers for both the encoder and the decoder on a generated dataset. The dataset consists of 2048 points sampled from one cycle of a sinusoidal wave, starting at a random phase in the cycle, i.e., $\sin([0,2\pi]+U(0,2\pi))$. With probability 0.5, this waveform is multiplied by an exponential function that smoothly increases from $e^{-1}$ to $e^{1}$; otherwise, it is multiplied by an exponential function that decreases from $e^{1}$ to $e^{-1}$. The autoencoder compressed the data to a low latent dimensionality of 2 (the number of independent features) to emphasize the differences between loss functions.
Figure 3(a) shows an example and its reconstruction after training with the conventional MSE loss, DSL only with $s=32$, and a combination of 1 MSE + 128 DSL. This may appear imbalanced, but was necessary for DSL to have an obvious effect on training, as the MSE loss value was usually about 150 times the DSL value. As expected, using MSE loss gives the closest reconstruction, but exhibits many direction changes compared to the original. Using only DSL, one may expect it to ideally go up or down linearly until the crest or trough of the wave, but this is not what happens. The reconstruction stays close to 0 while following the direction of the original curve, occasionally jumping back. This tendency arises from a combination of the low latent dimensionality and the properties of DSL. Outputs are encouraged to stay near 0 to avoid large penalties from mismatches and to allow the parameters to easily change the direction of the trend, as all it has to do is get the direction right. While the jump across 0 may appear large from the perspective of the absolute value of finite differences, they only cost one wrong sign in DSL.
Figure 3(b) shows the cumulative signs, which increase by 1 when the sign of the finite difference is positive, and decrease if it is negative. This can be thought of as the ideal output using only DSL. We expect constantly increasing or decreasing cumulative signs, which change directions at a sharp point, shown in the original. Using the combination of the loss functions, the reconstruction exhibits characteristics of both previous cases, showing a curve that resembles the shape of the original curve, with mostly the same trends, and a more sawtooth-like cumulative signs curve.
(a) values
(b) cumulative signs
Figure 3: Example originals and reconstructions of the generated sinusoidal wave dataset. The cumulative signs (3(b)) increase by 1 when the sign of the finite difference is positive, and decrease when it is negative. The combination of training with 1 MSE + 128 DSL results in reconstructions that have similar shape to the original and share movement directions.
This demonstration shows two extremes in a continuum of loss function choice. From the perspective of conventional distances, MSE does well and DSL does poorly, while from the perspective of behavior comparison, DSL does well while MSE does poorly. This suggests that in the typical case, where the preservation of trends is relevant, a mixture of loss functions would give the most desirable result. Of course, with sufficient resources (parameters, precision, etc.) either loss function encourages perfect reproduction of critical points. However, with limited resources, we must select different loss functions to emphasize different objectives.
4 Experiments
DSL can be used with MSE loss in most learnable compressor training processes. To assess its impact, we trained autoencoders on 1-dimensional arrays of stock prices, 2-dimensional images, and 3-dimensional physics simulation states. Key hyperparameters are in Appendix Table 1.
4.1 Stock prices
Stock prices are challenging to compress because price changes resemble Brownian motion (Samuelson, 2015). Certain trading strategies focus on the direction of price changes instead of the magnitude or absolute price (Bakhach et al., 2016; Aloud, 2015). We downloaded minute-level prices of NVDA from 2024-12-24 to 2025-01-22, totaling 15360 prices. We sampled subsequences of 64 minutes, normalized them to start at 0, and divided them by the overall standard deviation. We trained an MLP on this data with a latent dimensionality of 16.
Figure 4 shows an example reconstruction after training with different combinations of MSE and DSL. As expected, low latent dimensionality (given the Brownian-like data) forces the MSE-dominant models to smooth over the minute-level movements. Figure 5 shows three topological dissimilarity metrics after training: persistence diagram distance, functional distortion distance between merge trees, and pairwise linear correlation distance. When we use persistence diagram distance, we take the average of the order-2 Wasserstein distances among all homology orders up to $d-1$.
(a) normalized prices
(b) cumulative signs
(c) color bar
Figure 4: Original normalized stock prices vs reconstructions after training a simple autoencoder using MSE, DSL, or a combination thereof. While the MSE-dominant models follow the original closest, they smooth over the minute-level movements. The proportions of movement signs that are the same between the original and reconstructions are shown on the right side of Figure 4(b). Models trained on mixtures of MSE and DSL are more often correct in direction than those trained on pure MSE or pure DSL.
(a) persistence distance
(b) merge tree distance
(c) correlation distance
Figure 5: Topological dissimilarity metrics after training an MLP on NVDA stock prices. Squares show means while faint dots show individual examples. Error bars show one standard deviation away from the mean. Note the split horizontal axes; the left sides show a finer scale. The Brownian-like movement of minute-level prices is difficult to capture with high compression, although small proportions of DSL $\approx$ 0.08 help capture movement direction.
4.2 Images
The field of learnable data compressors has had the most study in the domain of image compression, particularly using variational autoencoders (VAEs), which are autoencoders whose latent representations are encouraged to be normally distributed. This allows generating new images by sampling from the normal latent distribution. We trained a quantization-aware ResNet-based VAE designed for image compression Duan et al. (2024), using different mixtures of MSE and DSL. We used Algorithm 2 on COCO 2017 Lin et al. (2014), the training dataset, to find a sharpness of 9.6.
(a) persistence distance
(b) merge tree distance
(c) correlation distance
(d) color bar
Figure 6: Topological metrics on a convolutional VAE trained on COCO 2017. Note the split axis on Figure 6(b). Each group on each reconstruction weight was trained on the same proportion of MSE and DSL, but the points are offset in the plots, and individual examples are omitted for clarity. Means and error bars are shown instead of individual examples. Reconstruction weight is a multiplier on the reconstruction loss as opposed to the KL divergence of the latent distribution from a normal distribution, a VAE optimization objective. Overall, 0.125 DSL proportion performs best.
Figure 6 shows topological metrics, and because we are compressing natural images, Appendix 6.6 shows visual fidelity metrics and reconstruction examples. Images were cropped to shape 3×64×643$\times$ 64$\times$ 643 × 64 × 64 on center when calculating the inefficient merge tree distance and correlation distance metrics.
4.3 Physical simulations
Shallow water simulations model fluid flow under the assumption that the water surface dimensions are much larger than its depth, which allows the simulations to avoid modeling vertical fluid velocity. These simulations output surface height and fluid velocities on horizontal axes. We used a simulation to generate data for evaluating DSL on a 3-dimensional dataset. Alternative datasets such as SDRBench Zhao et al. (2020) had only hundreds of examples or were best described as 2-dimensional with multiple channels (e.g. images). Our simulation initializes a shallow container of water with a bump in a random location. As time steps progress, the bump falls and the water moves around. We trained an autoencoder with two 3-dimensional convolutional layers for both the encoder and decoder, with $\mathrm{ReLU}$ between layers. Examples were shaped (3 channels, 8 snapshots, 64 length, 64 width). DSL sharpness was 16 and weight in the channel dimension was 0, treating each example as a tensor field with elements of (water height, velocityx, velocityy).
(a) persistence distance
(b) merge tree distance
(c) correlation distance
Figure 7: Topological measures on the 3D autoencoder trained on shallow water simulation states. Because VTK (Schroeder et al., 2006), the library we used for merge tree construction, supports a maximum of 3 dimensions, we evaluated the three channels separately. Models trained on $\approx$ 0.1 DSL perform the best overall.
Figure 7 shows topological metrics after training on mixtures of MSE and DSL. The best overall performance is achieved at $\approx$ 0.1 DSL proportion, suggesting that the most desirable mixture is nearby. Although the minimum correlation distance is achieved when using pure DSL, it results in obvious jumping-across-zero patterns that we observed in § 3.3. These patterns correspond to the mapping between the convolutional decoders and their inputs, where each input element corresponds to a $\times 4\times 4$ 4 block in the reconstruction (See Appendix 11 for visualizations). This suggests that the convolutional decoder layers have learned to output the correct water surface height gradients while the other layers may have focused on other features.
5 Limitations
As demonstrated in the stock prices experiment (§ 4.1), using DSL may not always have a significant benefit. Because learning behavior is sensitive to the balance between the DSL and traditional loss functions, due to their measuring fundamentally different quantities (e.g. squared errors versus counts of sign differences), a method for efficiently finding an optimal balance would facilitate its use. Finally, while DSL can be introduced into the training process of any typical autoencoder, we have not yet explored model architectures besides MLPs and CNNs. A task requiring a transformer-based autoencoder would bring additional insights into DSL’s effects.
6 Conclusion
We introduced directional sign loss (DSL) and investigated its properties and effects on training a variety of autoencoders on diverse datasets. We found that DSL, when used in combination with MSE loss, can serve as a proxy for common non-differentiable and computationally expensive topological feature descriptors, namely the Wasserstein distance between persistence diagrams, the functional distortion distance between merge trees, and the pairwise linear correlation distance, and can encourage autoencoders to preserve topological features such as critical points in their reconstructions.
Acknowledgments and Disclosure of Funding
Supported in part by NSF CISE Award 2217154.
Appendix
6.1 Making 0-dimensional persistence diagrams and taking the Wasserstein distance between them
Creating 0-dimensional persistence diagrams from tensors involves analyzing the connected components of sublevel sets of the scalar field defined by the tensor.
Input: The input is a tensor $T$ of arbitrary dimensionality $d$, representing a scalar field $f:X\to\mathbb{R}$, where $X$ is a $d$-dimensional grid.
6.1.1 Persistence diagram
To analyze the topology of the scalar field, we consider its sublevel sets $X_{t}=\{x\in X\mid f(x)\leq t\}$ As the threshold $t$ increases, the sublevel sets $X_{t}$ change, and we can track how their topology evolves. This process creates a filtration, a sequence of nested subspaces $X_{t_{1}}\subseteq X_{t_{2}}\subseteq\dots\subseteq X_{t_{n}}$.
To build a persistence diagram,
  1. 1.
    Flatten the tensor into a list of scalar values and their corresponding grid positions.
  2. 2.
    Sort the values in ascending order: $f_{1}\leq f_{2}\leq\dots\leq f_{m}$where $m$ is the total number of grid points.
  3. 3.
    Use a union-find data structure to keep track of connected components. Initialize each grid point as its own component.
  4. 4.
    For each point $x_{i}$ in order of increasing $f(x_{i})$:
    1. (a)
      If $x_{i}$ is a local minimum (i.e., all its neighbors in the grid have higher values), a new connected component is born. The definition of neighbors for each grid point is its orthogonally and diagonally adjacent elements. For a $d$-dimensional grid, each point has $3^{d}-1$ neighbors.
      • In 1D, each element has up to 2 neighbors.
      • In 2D, each pixel has up to 8 neighbors.
      • In 3D, each voxel has up to 26 neighbors.
    2. (b)
      If $x_{i}$ connects two existing components, the younger component dies and merges into the older one.
    3. (c)
      For each component, record its birth time (when it appears) and death time (when it merges with another component). A component that never dies is recorded as having a death time of +$\infty$+ ∞.
Time and memory complexity
The time complexity for computing a 0D persistence diagram for a tensor is:
  • Sorting: $O(\log m)$ using parallel sorting, where $m$ is the number of grid points.
  • Union-find operations: $O(m\alpha(m))$, where $\alpha(m)$ is the inverse Ackermann function (effectively constant).
For storing simplices, $O(m^{d})$ memory is necessary, where $d$ is the dimensionality of the tensor.
6.1.2 Example persistence diagram
Consider a 2D tensor $T=\begin{bmatrix}1&3&2\\ 4&1&5\\ 2&6&3\end{bmatrix}$
  1. 1.
    Flatten and sort the values: [1,1,2,2,3,3,4,5,6][1,1,2,2,3,3,4,5,6][ 1 , 1 , 2 , 2 , 3 , 3 , 4 , 5 , 6 ], keeping track of their locations.
  2. 2.
    Process the points.
    1. (a)
      At $t=1$: Two components are born (local minima at positions (1, 1) and (2, 2)).
    2. (b)
      At $t=2$: The components at (1, 3) and (3, 1) merge with existing components.
    3. (c)
      At $t=3$: The components at (1, 2) and (3, 3) merge with existing components.
    4. (d)
      At $t=4$: The component at (2, 1) merges with the component at (1, 1).
    5. (e)
      At $t=5$: The component at (2, 3) merges with the component at (2, 2).
    6. (f)
      At $t=6$: The component at (3, 2) merges with the component at (2, 2).
The persistence diagram will have points (1,2),(1,3),(1,4),(1,5),(1,6)(1,2),(1,3),(1,4),(1,5),(1,6)( 1 , 2 ) , ( 1 , 3 ) , ( 1 , 4 ) , ( 1 , 5 ) , ( 1 , 6 ).
6.1.3 Wasserstein distance
The Wasserstein distance (also known as the earth mover’s distance) between two persistence diagrams is a measure of the cost of transforming one diagram into the other.
The Wasserstein distance between two persistence diagrams $D_{1}$ and $D_{2}$ is defined as $\inf_{\gamma}(\sum_{(x,y)\in\gamma}\|x-y\|^{p})^{\frac{1}{p}}$ where $\gamma$ is a matching (bijection) between the points of $D_{1}$ and $D_{2}$, $\|x-y\|$ is the Euclidean distance between points $x$ and $y$, and $p$ is the order of the Wasserstein distance (typically $p=1$ or $p=2$).
To compute the Wasserstein distance between two persistence diagrams $D_{1}$ and $D_{2}$:
  1. 1.
    Add projections of all points in $D_{1}$ and $D_{2}$ onto the diagonal $b=d$ to ensure the two diagrams have the same number of points. Let $D_{1}^{\prime}$ and $D_{2}^{\prime}$ be the augmented diagrams.
  2. 2.
    Compute the pairwise Euclidean distances between all points in $D_{1}^{\prime}$ and $D_{2}^{\prime}$. This results in a cost matrix $C$, where $C_{ij}=\|x_{i}-y_{j}\|$ for $x_{i}\in D_{1}^{\prime}$ and $y_{j}\in D_{2}^{\prime}$.
  3. 3.
    Find the optimal matching $\gamma$ that minimizes the total cost $\sum_{(x,y)\in\gamma}\|x-y\|$.
  4. 4.
    Sum the costs of the optimal matching and take the $p$-th root.
Time and memory complexity
If $n$ is the number of pairs, the time complexity of computing the Wasserstein distance is dominated by solving the assignment problem: The cost matrix construction can be parallelized to $O(n)$ with infinite threads. However, the Hungarian algorithm is inherently sequential, and no known parallel algorithm can reduce its complexity below $O(n^{2})$. Thus, the overall time complexity is $O(n^{2})$.
The cost matrix takes $O(n^{2})$ space and the Hungarian algorithm requires $O(n^{2})$ space.
Thus, the total memory complexity is $O(n^{2})$.
6.2 Lower time and higher memory complexity than implemented
When calculating DSL, all $\Delta$ arrays for each dimension can be created in constant time, requiring $O(d\prod_{i=1}^{d}n_{i})$ memory to store. Then, all $\Delta$, containing a total of $d\prod_{i=1}^{d}(n_{i}-1)$ elements, are summed. Ignoring the subtraction of 111 for large $n_{i}$, this takes $O(\log_{2}(d\prod_{i=1}^{d}n_{i}))$ time, as elements must be summed two at a time. Taking the logarithm of the product, we have $O(\log_{2}d+\sum_{i=1}^{d}\log_{2}n_{i})$.
6.3 Loss function time and memory experiments with 3D inputs
Figure 8 shows time and memory usage with 3D cube-shaped input arrays.
(a) time 3D
(b) memory 3D
(c) legend
Figure 8: Time and peak memory usage from 3-dimensional array inputs using one AMD Ryzen 7 7800X3D CPU and one NVIDIA RTX 3090 GPU (24GB). Points are not shown if the time per example exceeds 1 second or GPU memory is exceeded. 3D inputs sizes show the number of elements per side of a cube. Persistence diagram distance was the 2-order Wasserstein distance between order-0 (the cheapest) persistence diagrams.
6.4 Hyperparameters used
Table 1 shows key hyperparameters used for all experiments.
Table 1: Hyperparameters used for experiments. Layer or filter counts and sizes are given as encoder; decoder. †For the Duan et al. [2024] model, they are: Number of filters: 3, 7, 6, 1, 6, 1, 4; 1, 8, 3, 2, 3, 4, 8. Filter sizes: 4, 7, 7, 2, 7, 2, 5, 2, 3; 1, 3, 5, 3, 7, 3, 7, 3, 7. Strides: 4, 1, 1, 2, 1, 2, 1, 2, 2; 1, 1, 1, 2, 1, 1, 2, 1, 1. Adam is [Kingma and Ba, 2017], He initialization is [He et al., 2015], and GELU is [Hendrycks and Gimpel, 2023].
Experiment Demo § 3.3 Stock prices § 4.1 Images § 4.2 Simulation § 4.3
Model architecture MLP MLP Duan et al. [2024] CNN
Layers or filters 2; 2 4; 4 64, 128; 128, 64
Layer or filter sizes 32; 32
2048, 256, 32; 32, 256, 2048
3, 3; 3, 3
Input size 2048 64 3 $\times$ 256 $\times$ 256 8 $\times$ 64 $\times$ 64
Latent dimension 2 16 512 4096
Activation function ReLU ReLU GELU ReLU
Optimizer Adam Adam Adam Adam
Scheduler none none cosine annealing 0.5 on 20 plateau
Initialization He He He He
Learning rate 0.001 0.0001 0.0002 0.0001
Dropout rate 0 0 0 0.1
Batches 20000 64000 50000 10000
Batch size 1024 64 32 64
Regularization 0 0 None 0
Sharpness 32 16 9.6 16
6.5 Choosing sharpness
Setting the sharpness hyperparameter $s$ can be tricky. We provide the following suggestion as a starting point: Set $s$ as a trainable parameter, and find the loss between random examples in the relevant dataset using an appropriate loss function, e.g. MSE loss for images. The MSE loss will tend toward some value. Then use stochastic gradient descent to minimize the difference between DSL and the reference loss function. Pseudocode is in Algorithm 2.
6.6 Natural image visual fidelity metrics and reconstruction example
(a) SSIM
(b) PSNR
(c) color bar
Figure 9: Visual fidelity metrics on a convolutional VAE trained on COCO 2017. Higher is better. Some subfigures have split axes. The structural similarity index measure (SSIM) and peak signal-to-noise ratio (PSNR) are measures of visual fidelity commonly used in audiovisual and signal processing work. Each group on each reconstruction weight was trained on the same proportion of MSE and DSL, but the points are offset in the plots, and individual examples are omitted for clarity.
Figure 10 shows an example image from the Kodak Lossless True Color Image Suite Eastman Kodak Company [1999] and its reconstructions, using the lowest reconstruction loss weight of 16, encouraging minimum bit rate. The only visible differences among the reconstructions are that using only DSL, the reconstruction tends to stay close to the middle of the dynamic range, resulting in a faded-looking reconstruction. This reflects the tendency of the reconstruction to stay near zero in § 3.3.
(a) Original
(b) MSE only
(c) DSL only
(d) 0.5 MSE + 0.5 DSL
Figure 10: An example image and its reconstructions from the validation dataset Eastman Kodak Company [1999]. Using only DSL (10(c)), the reconstruction tends to stay close to the middle of the dynamic range, resulting in a faded-looking reconstruction. We cannot see any differences between the others and the original, as reflected in Figure 9, where even the worst models in terms of visual fidelity (except pure DSL) perform acceptably.
Input: Dataset $\mathcal{D}$, reference loss $L_{\text{ref}}$, maximum steps $T$, stopping threshold $\epsilon$
Output: Optimized sharpness parameter $s$
Initialize $s$ as a trainable parameter;
Set optimizer $\mathcal{O}$;
for step $t\leftarrow 1$ to $T$ do
   Randomly sample a batch $\mathcal{B}\subset\mathcal{D}$;
   Split $\mathcal{B}$ into random subsets $\mathcal{B}_{1}$ and $\mathcal{B}_{2}$;
  Compute the reference loss:
$\ell_{\text{ref}}\leftarrow L_{\text{ref}}(\mathcal{B}_{1},\mathcal{B}_{2})$
  Compute the DSL loss using $s$:
$\ell_{\text{DSL}}\leftarrow\text{DSL}(\mathcal{B}_{1},\mathcal{B}_{2},s)$
  Compute the optimization loss:
$\ell_{\text{opt}}\leftarrow(\ell_{\text{DSL}}-\ell_{\text{ref}})^{2}$
  Backpropagate $\ell_{\text{opt}}$ to compute gradients of $s$;
   Update $s$ using $\mathcal{O}$ and reset gradients;
 if $\ell_{\text{opt}}<\epsilon$ then
    break;
return $s$
Algorithm 2 Find Sharpness
6.7 Shallow water reconstruction example
Figure 11 shows an example from the shallow water simulation dataset and its reconstructions using varying mixtures of MSE and DSL.
Figure 11: Original shallow water simulation snapshots and reconstructions. Column labels show loss function proportions used during training. Water surface height, velocityx, and velocityy are shown as red, green, and blue color channels.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract summarizes the paper.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations are discussed in the conclusion.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Theoretical claims are supported by experiments.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The main features are described in sufficient detail for reimplementation.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: Code is in supplementary material and will later be in a public repository.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [No]
  29. Justification: All details will not fit. Code is provided.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: Error bars are shown where appropriate.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Experiments dependent on hardware include hardware descriptions.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The work and anticipated effects conform to the Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: We do not anticipate any notable direct societal impact.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: We see no such threats.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: Original authors of assets are credited.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer:[Yes]
  64. Justification: New assets are described in the paper and included in supplementary material.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: The paper does not use crowdsourcing or experiment with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: The paper does not use crowdsourcing or experiment with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A] .
  79. Justification: The core method development of the work did not involve LLMs.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.